# Single-file dynamics with general charge measures

### v2

@@ 3 Dressing
where $n_{\pm}$ denote the number of particles crossing from right-to-left and vice-versa and $q_{c}^{\pm}\equiv\sum_{m_{\pm}=1}^{n_{\pm}}c_{m_{\pm}}$ is the total charge carried by the respective crossing particles. The single-file condition (8) implies that at most one of $n_{\pm}\neq 0$ which simplifies the conditional probability (16) to
@@
where $\nu=-{\rm sgn}({J}_{p})$. The conditional probability is then (up to a sign) the distribution of the sum of $|J_{p}|$ charges sampled from $\omega_{\mu}$. By the convolution theorem, the conditional distribution is therefore simply
@@
where we have introduced the Fourier transform of the charge measure
$\hat{\omega}_{\mu}(k)\equiv\int_{\mathcal{C}}{\rm d}\omega_{\mu}\,e^{-{\rm i}kc}.$ (19)
Dressing the particle distribution via factorization of the joint probability distribution (15) and the conditional distribution (18) can be interpreted as the action of a dressing operator $\mathfrak{D}_{\mathcal{P}}$
@@
Consider the Gaussian charge measure
$\omega_{\mu}^{\rm Gauss}(c)=\frac{e^{-(c-\mu)^{2}/{2\sigma_{\omega}^{2}}}}{\sqrt{2\pi\sigma_{\omega}^{2}}},$ (21)
with the Fourier transform
$\hat{\omega}_{\mu}^{\rm Gauss}(k)=e^{-\mu^{2}/2\sigma_{\omega}^{2}}e^{-(k\sigma_{\omega}+{\rm i}\mu/\sigma_{\omega})^{2}/2}.$ (22)
Computing the integral (18), we find the exact conditional probability
$\mathcal{P}^{\rm Gauss}_{c|p}({J}_{c}|{J}_{p})=\frac{e^{-(J_{c}-J_{p}\mu)^{2}/{2|J_{p}|\sigma_{\omega}^{2}}}}{\sqrt{2\pi|J_{p}|\sigma_{\omega}^{2}}}.$ (23)
3.1 Dressing the full-counting statistics
@@
In Appendix A the conjugation is computed explicitly, resulting in a straightforward substitution rule for exponentials of the particle counting field
@@
We note that $\hat{\omega}_{\mu_{\mp}}(\pm{\rm i}\lambda_{c})=\int_{\mathcal{C}}{\rm d}\omega_{\mu_{\mp}}e^{\pm c\lambda_{c}}$ are the charge full-counting statistics of a single particle going left-to-right or right-to-left respectively. Since charges are independent in the initial measure (6), the corresponding $n$-particle result factorizes in terms of the single-particle one. The substitution rule (26) therefore dresses the particle full-counting statistics by adjoining the left/right-going charge-full counting statistics with fixed particle number to the corresponding particle crossings.
@@
Consider the $\mathbb{Z}_{2}$ charge measure
@@
4 Dressing asymptotics
The dressing operator (26) gives the exact finite-time dressing of particle full-counting statistics. We now consider how the dressing stemming from the single-file condition (8) is manifested in asymptotic charge fluctuations on typical and large scales. We show that typical fluctuations depend only on the first two moments of the charge measure. Under the assumption of asymptotically Gaussian typical particle fluctuations, we derive typical charge fluctuations. We find that typical charge fluctuations in equilibrium ensembles with vanishing mean charge are anomalous (non-normal).
In practice, Gaussianity of typical fluctuations is commonly established as a consequence of a large deviation principle [3]. To proceed, we make the following assumptions about the particle full-counting statistics:
@@ 4 Dressing asymptotics
    Asymptotic values of scaled cumulants of the integrated particle current $s_{n}^{(p)}\equiv\lim_{t\to\infty}t^{-\alpha}\partial^{n}_{\lambda_{p}}\log G_{p}(\lambda_{p}|t)|_{\lambda_{p}=0}$ are finite, $|s_{n}^{(p)}|<\infty$, for all $n\in\mathbb{N}$ and the asymptotic value of the second cumulant (variance) is nonzero, $\sigma_{p}^{2}\equiv s_{2}^{(p)}\neq 0$.
Existence of $F_{p}(\lambda_{p})$ in assumption (I) can be reformulated as the exponential decay of asymptotic probabilities of large (of order $\mathcal{O}(t^{\alpha})$) fluctuations of the integrated particle current $J_{p}$
@@
where $I_{p}(j_{p})\equiv-\lim_{t\to\infty}t^{-\alpha}\log\mathcal{P}_{p}(J_{p}=j_{p}t^{\alpha}|t)$ is the particle rate function and $\asymp$ indicates asymptotic logarithmic equivalence. Since $F_{p}$ is assumed to be twice differentiable, the rate function $I_{p}$ is recovered by the Legendre-Fenchel transform [3]
@@
and is a twice differentiable strictly convex function on its domain $\mathcal{J}_{p}\equiv(j_{p}^{\rm min},j_{p}^{\rm max})$. We note that finiteness of scaled cumulants in (II) generally does not follow directly from the existence of the limit in (I), see [13] for a many-body system example. Instead, a sufficient condition is that the logarithm of the particle full-counting statistics $\log G_{p}(\lambda_{p}|t)$ is complex-analytic at the origin $\lambda_{p}=0$ for all $t$, see [9]. The assumption (II) should be understood as a regularity condition on the particle dynamics that ensures central limit behavior on the typical scale.
Taken together, assumption (I) and (II) imply that typical (of order $\mathcal{O}(t^{\alpha/2}$)) fluctuations of the integrated particle current $J_{p}$ around the average current (first cumulant) $\langle J_{p}(t)\rangle=s_{1}^{(p)}t^{\alpha}$ are asymptotically normally distributed
$\mathcal{P}^{\rm typ}_{p}(j_{p})\equiv\lim_{t\to\infty}t^{\alpha/2}\mathcal{P}_{p}(J_{p}=s_{1}^{(p)}t^{\alpha}+{j}_{p}t^{\alpha/2}|t)=\frac{e^{-{j}_{p}^{2}/2\sigma_{p}^{2}}}{\sqrt{2\pi\sigma_{p}^{2}}}.$ (32)
4.1 Typical charge fluctuations
We start by analyzing asymptotic typical charge fluctuations and show that they depend only on the mean and variance of the charge measure in addition to the particle variance.
We first consider a generic non-equilibrium initial ensemble with unequal initial particle densities $\rho_{+}\neq\rho_{-}$ and non-vanishing charge means $\mu_{\pm}\neq 0$. The difference of particle densities drives average particle and charge currents $s_{1}^{(i)}\neq 0$. To access typical fluctuations around these average values, we set $J_{i}=s_{1}^{(i)}t^{\alpha}+{j}_{i}t^{\alpha/2}$ and consider the joint typical distribution
$\mathcal{P}^{\rm typ}_{c,p}({j}_{c},{j}_{p})\equiv\lim_{t\to\infty}t^{\alpha}\mathcal{P}_{c,p}(J_{c}=s_{1}^{(c)}t^{\alpha}+{j}_{c}t^{\alpha/2},J_{p}=s_{1}^{(p)}t^{\alpha}+{j}_{p}t^{\alpha/2}).$ (33)
The factorized form of the joint typical distribution (15) at the typical scale then reads
$\mathcal{P}^{\rm typ}_{c,p}({j}_{c},{j}_{p})=\mathcal{P}^{\rm typ}_{c|p}({j}_{c}|{j}_{p})\mathcal{P}^{\rm typ}_{p}({j}_{p}),$ (34)
where we defined the typical conditional probability as
$\mathcal{P}^{\rm typ}_{c|p}({j}_{c}|{j}_{p})\equiv\lim_{t\to\infty}t^{\alpha/2}\mathcal{P}_{c|p}(J_{c}=s_{1}^{(c)}t^{\alpha}+{j}_{c}t^{\alpha/2}|J_{p}=s_{1}^{(p)}t^{\alpha}+{j}_{p}t^{\alpha/2}).$ (35)
As shown in Appendix B, the integral representation (18) allows us to relate the average values of charge and particle currents, giving the expected result
@@ 4.1 Typical charge fluctuations
where $\overline{\nu}=-{\rm sgn}(s_{1}^{(p)})$. Computing the typical conditional probability we find
$\mathcal{P}^{\rm typ}_{c|p}({j}_{c}|{j}_{p})=\frac{e^{-({j}_{c}-{j}_{p}\mu_{\overline{\nu}})^{2}/2|s_{1}^{(p)}|\sigma_{\omega}^{2}}}{\sqrt{2\pi|s_{1}^{(p)}|\sigma_{\omega}^{2}}}.$ (37)
The typical charge distribution is obtained by marginalizing the typical joint distribution (34) (for $\mu_{\overline{\nu}}\neq 0$), yielding a Gaussian distribution
$\mathcal{P}^{\rm typ}_{c}({j}_{c})=\int_{\mathbb{R}}{\rm d}j_{p}\,\mathcal{P}^{\rm typ}_{c|p}({j}_{c}|{j}_{p})\mathcal{P}^{\rm typ}_{p}({j}_{p})=\frac{e^{-j_{c}^{2}/2\sigma_{c}^{2}}}{\sqrt{2\pi\sigma_{c}^{2}}},$ (38)
with charge variance
$\sigma_{c}^{2}=\mu^{2}_{\overline{\nu}}\sigma_{p}^{2}+|s_{1}^{(p)}|\sigma^{2}_{\omega}.$ (39)
Eq. (39) shows that the charge variance decomposes into a contribution stemming from typical particle fluctuations with variance $\sigma_{p}^{2}$ and average charge $\mu_{\overline{\nu}}$ and a contribution from the variance of the charge measure $\sigma_{\omega}^{2}$ carried by the average particle current $|s_{1}^{(p)}|$.
@@
$\mathcal{P}^{\rm typ,\,eq}_{c|p}({j}_{c}|{j}_{p})\equiv\lim_{t\to\infty}t^{\alpha/4}\mathcal{P}_{c|p}(J_{c}={j}_{c}t^{\alpha/4}|J_{p}={j}_{p}t^{\alpha/2}).$ (40)
As shown in Appendix B, Eq. (18) now gives
Remarkably, the resulting typical charge distribution $\mathcal{P}^{\rm typ,eq}_{c}(j_{c})\equiv\lim_{t\to\infty}t^{\alpha/4}\mathcal{P}_{c}(J_{c}={j}_{c}t^{\alpha/4}|t)=\int_{\mathbb{R}}{\rm d}j_{p}\,\mathcal{P}^{\rm typ,\,eq}_{c|p}({j}_{c}|{j}_{p})\mathcal{P}^{\rm typ}_{p}({j}_{p})$ is not Gaussian,
$\mathcal{P}^{\rm typ,eq}_{c}({j}_{c})=\int_{\mathbb{R}}\frac{{\rm d}j_{p}e^{-j_{c}^{2}/2|j_{p}|\sigma_{\omega}^{2}-j_{p}^{2}/2\sigma_{p}^{2}}}{2\pi\sigma_{p}\sigma_{\omega}|j_{p}|^{1/2}}=\frac{M_{1/4}(2|j_{c}|/\sigma_{\omega}\sigma_{p}^{1/2})}{\sigma_{\omega}\sigma_{p}^{1/2}},$ (42)
but is instead given by the M-Wright distribution $M_{1/4}$ [16]. This distribution has been shown to govern charge fluctuations in inert $\mathbb{Z}_{2}$-charged single-file systems [13, 21] and Eq. (42) extends the result to arbitrary charge measures with finite variance.
4.2 Large charge fluctuations
Typical charge fluctuations depend only on the first two moments of the charge measure, whereas large fluctuations are sensitive to the full structure of the measure. In the following analysis it will be convenient that the Fourier transform of the measure $\hat{\omega}_{\mu}(k)$ is an entire function. By the Paley-Wiener theorem [27], a sufficient assumption is that the charge measure decays faster than exponentially on the real line
$\omega_{\mu}(c)<ae^{-b|c|}\quad{\rm as}\enspace|c|\to\infty,\quad\textrm{for fixed $a$ and all }b\in\mathbb{R^{+}}.$ (43)
The joint charge-particle rate function $I_{c,p}(j_{c},j_{p})=-\lim_{t\to\infty}t^{-\alpha}\log\mathcal{P}_{c,p}(J_{c}=j_{c}t^{\alpha},J_{p}=j_{p}t^{\alpha})$ is obtained by taking the logarithm of the factorized probability (15)

### v3

@@ III Dressing
where $n_{\pm}$ denote the number of particles crossing from right-to-left and vice-versa and $q_{c}^{\pm}\equiv\sum_{m=1}^{n_{\pm}}c_{m}$ is the total charge carried by the respective crossing particles. Some comments on Eq. (13) are in order. Since inert charges do not influence particle dynamics, the conditional probability is obtained by first summing over all combinations of particles crossing the origin (first sum and $\delta$-function in Eq. (13)) and then averaging over possible charge configurations of these particles (second $\delta$-function in Eq. (13)). Given that charges are inert and not correlated in the initial measure (7), this amounts to averaging over independent charge distributions (factorized charge measure in Eq. (13)).
For a single-file dynamics (8), either all particle cross from left to right or vice versa, so that at most one of $n_{\pm}\neq 0$ which simplifies the conditional probability (13)
@@
where $\nu=-{\rm sgn}({J}_{p})$. The conditional probability is then (up to a sign) the distribution of the sum of $|J_{p}|$ charges sampled from $\omega_{\mu}$. By the convolution theorem, the conditional distribution reads
@@
where we have introduced the Fourier transform of the charge measure $\hat{\omega}_{\mu}(k)\equiv\int_{\mathcal{C}}{\rm d}\omega_{\mu}\,e^{-{\rm i}kc}.$ We denote the dressing of the particle distribution via factorization of the joint probability distribution and the conditional distribution (15) as the action of a dressing operator $\mathfrak{D}_{\mathcal{P}}$
@@
An exactly solvable example of dressing is given by a Gaussian charge measure
$\omega_{\mu}^{\rm Gauss}(c)=\mathcal{N}_{c}(\mu,\sigma_{\omega}^{2}),$ (17)
where $\mathcal{N}_{x}(\overline{x},\sigma^{2})\equiv e^{-(x-\overline{x})^{2}/2\sigma^{2}}/\sqrt{2\pi\sigma^{2}}$, owing to additivity of variances of independent Gaussian variables. The Fourier transform of Eq. (17) reads
$\hat{\omega}_{\mu}^{\rm Gauss}(k)=2\pi\mathcal{N}_{\mu}(0,\sigma_{\omega}^{2})\mathcal{N}_{k}(-{\rm i}\mu\sigma_{\omega}^{-2},\sigma_{\omega}^{-2}).$ (18)
Eq. (15) reduces to a Gaussian integral which gives the exact conditional probability
$\mathcal{P}^{\rm Gauss}_{c|p}({J}_{c}|{J}_{p})=\mathcal{N}_{J_{c}}(J_{p}\mu,|J_{p}|\sigma_{\omega}^{2}).$ (19)
Note that while Eq. (15) gives the exact conditional probability for an arbitrary charge measure, the integral generally cannot be evaluated in closed form as for a Gaussian measure and must instead be evaluated asymptotically, see Section IV.
III.1 Dressing the full-counting statistics
@@
The conjugation is computed in Appendix A, resulting in a straightforward substitution rule for exponentials of the particle counting field
@@
We note that $\hat{\omega}_{\mu_{\mp}}(\pm{\rm i}\lambda_{c})=\int_{\mathcal{C}}{\rm d}\omega_{\mu_{\mp}}e^{\pm c\lambda_{c}}$ are the charge full-counting statistics of a single particle going left-to-right or right-to-left respectively. Since charges are uncorrelated in the initial measure (6), the corresponding $n$-particle result factorizes in terms of the single-particle one. The substitution rule (22) therefore dresses the particle full-counting statistics by adjoining the left/right-going charge full-counting statistics with fixed particle number to the corresponding particle crossings.
@@
The $\mathbb{Z}_{2}$ charge measure reads
@@
IV Dressing asymptotics
The dressing operator (22) gives the exact finite-time dressing of particle full-counting statistics. We now consider how the dressing manifests in time-asymptotic charge fluctuations on typical and large scales which we specify below. We show that typical fluctuations depend only on the first two moments of the charge measure. Assuming asymptotically Gaussian typical particle fluctuations, we demonstrate that typical charge fluctuations in equilibrium ensembles with vanishing mean charge are anomalous (non-normal).
Gaussianity of typical fluctuations is commonly established as a consequence of a large deviation principle [3]. To proceed, we make the following assumptions about the particle full-counting statistics:
@@ IV Dressing asymptotics
    Asymptotic values of scaled cumulants $s_{n}^{(p)}\equiv\lim_{t\to\infty}t^{-\alpha}\partial^{n}_{\lambda_{p}}\log G_{p}(\lambda_{p}|t)|_{\lambda_{p}=0}$ are finite, $|s_{n}^{(p)}|<\infty$, for all $n\in\mathbb{N}$ and the asymptotic value of the second cumulant is nonzero, $\sigma_{p}^{2}\equiv s_{2}^{(p)}\neq 0$.
The asymptotic exponential decay of large (of order $\mathcal{O}(t^{\alpha})$) fluctuations of the integrated particle current $J_{p}$ is governed by a large deviation principle [3]
@@
where $I_{p}(j_{p})\equiv-\lim_{t\to\infty}t^{-\alpha}\log\mathcal{P}_{p}(J_{p}=j_{p}t^{\alpha}|t)$ is the particle rate function and $\asymp$ indicates asymptotic logarithmic equivalence. The twice differentiable $F_{p}(\lambda_{p})$ in assumption (i) is related to the rate function $I_{p}$ is related by the Legendre-Fenchel transform
@@
which shows that $I_{p}$ is a twice differentiable strictly convex function on its domain $\mathcal{J}_{p}\equiv(j_{p}^{\rm min},j_{p}^{\rm max})$.
While large fluctuations become exponentially unlikely at large times, typical (of order $\mathcal{O}(t^{\alpha/2})$) fluctuations have a finite asymptotic distribution. Noting that the average integrated particle current is given asymptotically by the first cumulant $\langle J_{p}(t)\rangle\simeq s_{1}^{(p)}t^{\alpha}$ we denote $J_{p}[j_{p}]=s_{1}^{(p)}t^{\alpha}+{j}_{p}t^{\alpha/2}$. The asymptotic typical distribution is then given by
$\mathcal{P}^{\rm typ}_{p}(j_{p})\equiv\lim_{t\to\infty}t^{\alpha/2}\mathcal{P}_{p}(J_{p}=J_{p}[j_{p}]|t).$ (28)
Taken together, assumption (i) and (ii) ensure that the typical distribution of $j_{p}$ is asymptotically normal
$\mathcal{P}^{\rm typ}_{p}(j_{p})=\mathcal{N}_{j_{p}}(0,\sigma_{p}^{2}).$ (29)
We note that finiteness of scaled cumulants in (ii) and the resulting asymptotic normality generally does not follow directly from the existence of the limit in (i). Instead, a sufficient condition is that the logarithm of the particle full-counting statistics $\log G_{p}(\lambda_{p}|t)$ is complex-analytic at the origin $\lambda_{p}=0$ for all $t$, see [9]. The assumption (ii) should be understood as a regularity condition on the particle dynamics that ensures central limit behavior on the typical scale set by the variance.
IV.1 Typical charge fluctuations
Consider a non-equilibrium initial ensemble (7) with generic particle densities $\rho_{+}\neq\rho_{-}$ and charge means $\mu_{\pm}\neq 0$. The difference of particle densities drives average particle and charge currents $s_{1}^{(i)}\neq 0$. To access typical fluctuations around the averages, we set $J_{i}[j_{i}]=s_{1}^{(i)}t^{\alpha}+{j}_{i}t^{\alpha/2}$ and consider the joint typical distribution $\mathcal{P}^{\rm typ}_{c,p}({j}_{c},{j}_{p})\equiv\lim_{t\to\infty}t^{\alpha}\mathcal{P}_{c,p}(J_{c}[j_{c}],J_{p}[j_{p}]|t).$ The factorized form of the joint typical distribution at the typical scale then reads $\mathcal{P}^{\rm typ}_{c,p}({j}_{c},{j}_{p})=\mathcal{P}^{\rm typ}_{c|p}({j}_{c}|{j}_{p})\mathcal{P}^{\rm typ}_{p}({j}_{p})$, where the typical conditional probability is $\mathcal{P}^{\rm typ}_{c|p}({j}_{c}|{j}_{p})\equiv\lim_{t\to\infty}t^{\alpha/2}\mathcal{P}_{c|p}(J_{c}[j_{c}]|J_{p}[j_{p}]).$ As shown in Appendix B.1, the integral representation (15) allows us to relate the average values of charge and particle currents, giving the expected result
@@ IV.1 Typical charge fluctuations
where $\overline{\nu}=-{\rm sgn}(s_{1}^{(p)})$ with Gaussian fluctuations around the mean value, $\mathcal{P}^{\rm typ}_{c|p}({j}_{c}|{j}_{p})=\mathcal{N}_{j_{c}}(j_{p}\mu_{\overline{\nu}},|s_{1}^{(p)}|\sigma_{\omega}^{2}).$ The typical charge distribution is obtained by marginalization, yielding a Gaussian distribution
$\mathcal{P}^{\rm typ}_{c}({j}_{c})=\mathcal{N}_{j_{c}}\left(0,\sigma_{c}^{2}\right),$ (31)
with charge variance $\sigma_{c}^{2}=\mu^{2}_{\overline{\nu}}\sigma_{p}^{2}+|s_{1}^{(p)}|\sigma^{2}_{\omega}.$ The charge variance decomposes into a contribution stemming from typical particle fluctuations with variance $\sigma_{p}^{2}$ and average charge $\mu_{\overline{\nu}}$ and a contribution from the charge measure variance $\sigma_{\omega}^{2}$ carried by the average particle current $|s_{1}^{(p)}|$.
@@
Figure 2: Typical distribution of transported charge $\mathcal{P}_{c}^{\rm typ,\,eq}$ (32) with $\sigma=1$ (full blue curves) in an equilibrium ensemble with vanishing mean charge in logarithmic (main figure) and linear (inset) scale. The distribution is strongly non-Gaussian, decaying asymptotically as $\mathcal{P}_{c}^{\rm typ,\,eq}(\sigma j_{c}/\sqrt{2})\simeq 2|j_{c}|^{-1/3}e^{-\tfrac{3}{4}|j_{c}|^{4/3}}/\sqrt{3\pi}$ (dashed red curves) and featuring a cusp at $j_{c}=0$.
Remarkably, the resulting typical charge distribution $\mathcal{P}^{\rm typ,eq}_{c}(j_{c})=\int_{\mathbb{R}}{\rm d}j_{p}\,\mathcal{P}^{\rm typ,\,eq}_{c|p}({j}_{c}|{j}_{p})\mathcal{P}^{\rm typ}_{p}({j}_{p})$ is non-Gaussian but is instead given by an M-Wright distribution, see Ref. [16], Appendix C and Figure 2
$\mathcal{P}^{\rm typ,eq}_{c}({j}_{c})=\sigma^{-1}M_{1/4}(2|j_{c}|/\sigma),$ (32)
where $M_{1/4}(|x|)=\int_{\mathbb{R}}\frac{{\rm d}y}{2\pi}|y|^{-1/2}e^{-x^{2}/4|y|-y^{2}/4}$ and $\sigma^{2}\equiv\sqrt{2}\sigma_{p}\sigma_{\omega}^{2}$. Compared to a Gaussian distribution, the anomalous distribution (32) has heavier tails with the leading asymptotic term of the M-Wright function given by Laplace’s method as $M_{1/4}(\sqrt{2}(2x)^{3})\simeq\tfrac{e^{-12x^{4}}}{\sqrt{3\pi x^{2}}}$ for $|x|\to\infty$. M-Wright functions of general order occur as fundamental solutions in stochastic processes described by fractional diffusion equations [15, 16, 18, 17]. In the realm of fluctuations in deterministic systems, the distribution (32) has already been found to govern charge fluctuations in inert $\mathbb{Z}_{2}$ charged single-file systems [13, 21]. Unlike in fractional diffusion equations, the order of the M-Wright function in single-file systems is fixed to $1/4$ independently of the time-scale of underlying dynamics. Our result shows that anomalous equilibrium typical fluctuations are independent of the detailed structure of the charge measure, requiring only a finite variance and the single-file property (8).
IV.2 Large charge fluctuations
Typical charge fluctuations depend only on the first two moments of the charge measure, whereas large fluctuations are sensitive to the full structure of the measure. In the following it will be convenient that the Fourier transform of the measure $\hat{\omega}_{\mu}(k)$ is an entire function. By the Paley-Wiener theorem [27], a sufficient assumption is that the charge measure decays faster than exponentially on the real line, $\log(\omega_{\mu}(c))/|c|\rightarrow-\infty$ as $|c|\to\infty$.
The joint charge-particle rate function $I_{c,p}(j_{c},j_{p})=-\lim_{t\to\infty}t^{-\alpha}\log\mathcal{P}_{c,p}(J_{c}=j_{c}t^{\alpha},J_{p}=j_{p}t^{\alpha})$ is obtained by taking the logarithm of the factorized probability